what us the molarity of "11.2"of H2O2

Grade:10

1 Answers

ratna teja honey
10 years ago

Now let us say we have a 1 M solution of H2O2. What is its volume strength and what is its normality?

2H2O2 --> 2H2O + O2


The oxidation number of oxygen changes from -2 to 0, so the n factor of hydrogen peroxide is 2. Hence its normality is 2 x molarity. So 1 M solution of hydrogen peroxide = 2N.

Now, what is the volume strength? 

Now 1 litre of H2O2 has 1 M of it.  This will produce 1/2 M of oxygen = 22.4/2 = 11.2 litres of oxygen since 1 mole of an ideal gas at STP occupies 22.4 litres of volume.

Hence, 1 litre of this particular H2O2 solution has a volume strength of 11.2 since 1 litre of this H2O2 solution produces 11.2 litres of oxygen on decomposition.